# Revisiting the Galactic Winds in M82 I: the recent starburst and launch of outflow in simulations

### v2

@@ 2.3 Star formation
Previous simulations of M82’s galactic winds often rely on prescribed star formation histories, such as instantaneous starbursts or constant star formation rates. Additionally, SN feedback has been injected into a thin disk plane (Strickland & Stevens 2000; Cooper et al. 2008), or a central 1 kpc spherical region based on assumed star cluster distributions (e.g.,Melioli et al. 2013;Schneider et al. 2020;Schneider & Mao 2024). A key goal of our work is to more self-consistently model the recent starburst in M82’s nuclear region and its associated feedback. To this end, both the star formation process and how they couple with the surrounding ISM are solved explicitly in our simulation.
Newly born stars are embedded in molecular clouds and are usually surrounded by overdense ISM. Recent simulations of dwarf galaxies, which resolve star formation down to the scale of molecular clouds, have employed various approaches to capture star formation process. For dwarf galaxies with masses of $10^{7}-10^{8}\rm{M_{\odot}}$ and relatively low SFR, a star-by-star formation recipe (Gutcke et al. 2021), is effective when sub-parsec resolution is achieved. However, for galaxy-scale simulations with coarser resolution (several parsecs or larger), a sink particle approach is more practical. This approach involves creating sink particles with masses exceeding that of a single star in high-density regions, allowing them to accrete mass from surrounding gas cells until they meet the criteria for spawning star particles, each representing a cluster of stars (e.g., Bate et al. 1995; Krumholz et al. 2004; Jappsen et al. 2005; Federrath et al. 2010b; Gong & Ostriker 2012; Gatto et al. 2016; Howard et al. 2016). We adopt a sink particle model based on the methods of Federrath et al. (2010b) and Howard et al. (2016) to solve the star formation process in our work..
@@
These criteria are commonly used for sink particle implementations in mesh-based fluid dynamic solvers (e.g. Krumholz et al. 2004; Federrath et al. 2010b; Gong & Ostriker 2012). The criterion requiring the cell to be gravitationally bound has also been adopted in many studies. However, some studies indicate that stars can form in areas that are not gravitationally bound, such as the fast-expanding shell of an SNR bubble or a wind bubble (Cantat-Gaudin et al. 2018; Deharveng et al. 2009) or high-velocity fragments in the outflow area (Maiolino et al. 2017). To account for these scenarios, we relax the gravitational bound criterion in our simulations. The following subsections detail these criteria and the star formation process from sink particles.
@@ 2.3.2 Distance check
Even if gas clouds near a newly formed star cluster meet the density threshold for star formation, they can be disrupted by photoionization heating and radiative pressure from massive stars. In M82, star clusters are often embedded in molecular clouds with optical depths of up to $\tau\sim 10$. In such environments, radiative pressure may be comparable to supernovae in terms of momentum injection (Agertz et al. 2012). Combined with stellar winds, these early feedback processes can suppress star formation in neighboring regions. To account for this, we assign each sink particle a control volume that should not overlap with those of existing sink or star particles. The side length of these cubic control volumes is set to three times the grid cell size.
@@ 2.3.3 Convergent flow
To confirm that a candidate cell is undergoing free-fall collapse, we check for convergent gas flow. This requires that the divergence of the velocity field , $\nabla\cdot v<0$, is negative in the central cell of the control volume (Gong & Ostriker 2012).
@@ 2.3.5 Gas accretion and star formation
Several methods exist for calculating gas accretion onto sink particles. Krumholz et al. (2004) categorize accretion into three scenarios based on the relative importance of pressure and gravity, as determined by the particle’s Bondi-Hoyle radius, and employ different accretion models accordingly. Gong & Ostriker (2012) introduce ‘ghost’ zones around particles and calculate accretion rates based on the flux at the zone boundaries. In our work, we adopt a simpler approach similar to Federrath et al. (2010b). Once a sink particle forms, gas accretion begins immediately. Grid cells within the control volume are continuously monitored, and any gas exceeding the density threshold $\rho_{th}$ is directly accreted by the sink particle. The mass accretion rate of a sink particle at each time step is given by:
@@
In our simulation, the radius of the control volume of a sink particle, $R_{cv}$, is $\sim 1.5$ times the size of the nearby grid. Since sink particles can form in regions with varying resolutions, their control volumes and the corresponding Jeans masses $M_{J}$ can differ. A fixed mass threshold, $M_{cnv}$, could lead to inconsistent distribution of stellar mass fraction, i.e. the ratio of stellar mass to the total mass in a star particle, across regions with different resolutions. To ensure a consistent distribution of the stellar mass fraction, we adopt a variable $M_{cnv}$. Based on the assumption that a) the total accreted mass is proportional to Jeans mass, b) sink particles of different sizes have a constant temperature, and c) the stellar mass fraction is independent of the particle mass, we derive $M_{cnv}\propto M_{J}\propto R_{cv}$. In our simulation, we adopt $M_{cnv}=5.0\times 10^{4}\,\rm{M_{\odot}}$ for the finest grids with $dx=4.0pc$, $M_{cnv}=1.0\times 10^{5}\,\rm{M_{\odot}}$ for $dx=8.0pc$, etc.
@@ 2.4 Feedback
Stellar feedback serves as the main driving force for the launch of the galactic outflow. Traditionally, the primary feedback mechanism considered in similar studies is the core-collapse SNe, which has been proven capable of launching galaxy-scale outflow in M82 (e.g. Strickland & Stevens 2000; Cooper et al. 2008; Schneider et al. 2020). However, recent studies show that radiation feedback and stellar wind may play crucial roles in manipulating the effect of SNe feedback (Dale et al. 2014; Chevance et al. 2020). These early feedback processes may significantly lower the SFR and prevent the clustering of star formation. Some previous work suggests that efficient X-ray and extreme ultraviolet (EUV) radiation are capable of removing important coolants through photoionization, leading directly to the reduction of the accretion rate and SFR (e.g. Cantalupo 2009). Furthermore, stellar wind can create cavities and channels of ionized gas before the first SNe, mitigating the resistance an expanding SNR may meet. In this study, we consider three major feedback processes: radiation, stellar wind, and core-collapse SNe. We will examine how these mechanisms affect the starburst activity and the launch of outflow in M82. The effects of Type Ia SNe are ignored because our simulations last for 30 Myr, in which period type Ia SNe should have a limited effect in contrast to the core-collapse SNe. In our simulations, the stellar mass of a single star particle is equivalent to a medium-sized star cluster. Therefore, it can continuously produce stellar wind and radiation feedback and it will launch several individual SNe throughout its lifetime.
@@ 2.4.1 Gas return from sink particle
Most previous studies that adopted a sink particle-based star formation scheme have not included any returning mechanism for the remaining gas in sink particles. Some of these studies may assume that all the gas in sink particles will fully convert to stars (e.g. Wang et al. 2010; Padoan & Nordlund 2011; Gong & Ostriker 2012; Kim et al. 2018), which is valid only if the resolution can reach 0.1 parsec or higher. However, this treatment may overestimate the star formation efficiency in simulations with a resolution coarser than the size of the cores of GMCs. Meanwhile, other simulation studies on the evolution of GMC and ultra-faint dwarf galaxies believe that most unresolved gas in sink particles remains gravitationally bound to the cluster formed in GMC over a long time scale (Dale et al. 2014; Howard et al. 2016). Therefore, gas is stored in sink particles in these studies. However, in simulations with a relatively coarser resolution or with a long simulation time, such a treatment would cause a great fraction of the gas to be locked up in particles (Truelove et al. 1997; Hu et al. 2017), leading to an underestimation of the star formation efficiency. Overall, both approaches mentioned above have their own limitations. Recent studies indicate that early pre-supernovae feedback processes (i.e. stellar wind, radiation feedback) are capable of destroying the cloud and accelerating a great portion of the remaining gas to a gravitationally unbound state within 2 $\sim$ 15 Myr, in which the returning fraction may depend on the properties of the gas clumps (e.g., Kim et al. 2018; Li et al. 2019).
In our simulation, we use a simplified model to account for the dispersal of gas remaining in the sink/star particles. In the default run, we assume that all the remaining gas in the particles can eventually escape from their hosts. For comparison, we have also adopted two different setups in other runs, one with a gas return fraction of $50\%$, the other with the gas return fully disabled. The return of the remaining gas in a particle begins 2.0 Myr after star formation, based on previous simulations (e.g. Dale et al. 2014; Kim et al. 2018; Li et al. 2019; Fukushima & Yajima 2021). Moreover, the gas stored in a particle will return to the surrounding grid cells as an exponential function of time on a time scale $\tau_{rtn}$
$\dot{M}_{gas}=-\frac{M_{gas}}{\tau_{rtn}}.$ (10)
Here in our simulation, a time scale of $\tau_{rtn}=5.0Myr$ is adopted. At each time step, gas leaked from a particle is injected into grid cells within the particle’s control volume with an initial velocity equal to that of the host particle, and subsequently accelerated by stellar winds and radiation pressure applied to the same region.
@@ 2.4.2 Radiation
Multiple researches have shown that radiation pressure and photoionization dominate over gas pressure in the central region of molecular clouds, acting as the main mechanisms for cloud destruction and disrupting star formation (Ceverino et al. 2013; Sales et al. 2013). More recently, radiation feedback has been proven capable of removing gas from the center of a molecular cloud and carving channels into the outer region, allowing supernova remnants to pass through without much resistance (Dale et al. 2014; Fujii et al. 2021). Emerick et al. (2018) shows that these two mechanisms can play a catalytic role in the launch of galactic outflow in a low-mass dwarf galaxy. To investigate the effect of radiation, we incorporate both radiation pressure and radiation heating feedback in our simulation. Note that photoionization is not explicitly included in our simulation. Instead, we employ photoelectric heating in our cooling/heating source term for the following reasons: 1) resolving the detailed chemical evolution is not the main intention of this work, and 2) most of the gas in M82 remains neutral throughout the starburst, while ionized gas in the dense area will recombine quickly. From a statistical point of view, photoionization can be treated as a heating term. In our simulation, only UV radiation is considered, as it carries the most energy and momentum at the early age of a starburst galaxy. We use two separate modules to account for the effect of the photoheating and radiation pressure separately, both of which will be described in this section.
@@
For photoelectric heating, we use equation 1 in Wolfire et al. (1995) to calculate the heating rate. This method has been implemented into the GRACKLE library and integrated with heating and cooling routines. To compute photoelectric heating, the intensity of the interstellar radiation field must first be determined. To this end, we use an approach different from the optically thin assumption, considering that most parts of the M82’s gas disk, especially the star forming clumps and SNR bubble shells, are opaque to UV radiation (Murray et al. 2009). Thompson et al. (2008) suggested that the characteristic opacity of UV radiation is $\kappa_{UV}\sim 10^{3}\,\rm{cm}^{2}\cdot\rm{g}^{-1}$ in the neutral gas. The two-temperature Planck mean opacity calculated from OPTAB (Hirose et al. 2021) also shows that for a typical OB star with an effective surface temperature of more than $10^{5}$ K surrounded by ambient gas with a temperature ranging from 10 to 20000 K, the opacity can be as high as $10^{2}\sim 10^{3}\,\rm{cm}^{2}\cdot\rm{g}^{-1}$. In such an environment, the inverse-square law could overestimate the intensity of the UV radiation field by a few orders of magnitude in a region far away from the UV source. To correct for this effect, we instead use a modified inverse-square law to estimate the intensity of the radiation field. More specifically, the UV flux distribution is obtained by solving the point-source divergence equation with absorption under spherical symmetry as follows.
@@
where $L$ is the length of the transmission path of the light. The IR optical depth $\tau_{IR}$ in M82’s star cluster is assumed to be $10\sim 100$ based on the observed densities (Agertz et al. 2012), therefore IR may be the dominant term in radiation pressure. In the calculation of optical depth, we use a similar approach to that in the calculation of self-shielding in GRACKLE (Smith et al. 2016), we use the local Jeans length $R_{J}$ as the radius of the host clump within which the grid cell resides, therefore
@@
$\frac{d\textbf{P}}{dt}=-\nabla P_{Radiation}$ (19)
@@ 2.4.3 Stellar Wind
In addition to radiation and SNe, the stellar wind is a non-negligible source of feedback, which could be comparable to that of SNe under certain circumstances (Agertz et al. 2012). We adopt a metallicity-dependent model proposed by Vink & Sander (2021) to calculate the stellar wind’s mass loss rate and final velocity.
Table 2: List of all the simulations performed, the first and second columns are the name and mass of gas disc, respectively. The third, fourth and fifth column indicate the setting of supernovae, stellar winds and radiation feedback, respectively. The sixth and seventh columns indicate the gas return fraction and initial gas metallicity respectively (see section 2.6 for more details).
@@
GL00 $3.0\times 10^{9}\rm{M_{\odot}}$ vSN(S16) Yes Yes $0\%$ $0.02Z_{\odot}$

### v4

@@ 2.3 Star formation
Previous simulations of galactic winds in M82-like galaxies often rely on prescribed star formation histories, such as instantaneous starbursts or constant star formation rates. Additionally, SN feedback has been injected into a thin disk plane (Strickland and Stevens 2000; Cooper et al. 2008), or a central 1 kpc spherical region based on assumed star cluster distributions (e.g.,Melioli et al. 2013;Schneider et al. 2020;Schneider and Mao 2024). A key goal of our work is to more self-consistently model the recent starburst in M82’s nuclear region and its associated feedback. To this end, both the star formation process and how they couple with the surrounding ISM are solved explicitly in our simulation.
Newly born stars are embedded in molecular clouds and are usually surrounded by overdense ISM. Recent simulations of dwarf galaxies, which resolve star formation down to the scale of molecular clouds, have employed various approaches to capture star formation process. For dwarf galaxies with masses of $10^{7}-10^{8}\rm{M_{\odot}}$ and relatively low SFR, a star-by-star formation recipe (Gutcke et al. 2021), is effective when sub-parsec resolution is achieved. However, for galaxy-scale simulations with coarser resolution (several parsecs or larger), a sink particle approach is more practical. This approach involves creating sink particles with masses exceeding that of a single star in high-density regions, allowing them to accrete mass from surrounding gas cells until they meet the criteria for spawning star particles, each representing a cluster of stars (e.g., Bate et al. 1995; Krumholz et al. 2004; Jappsen et al. 2005; Federrath et al. 2010b; Gong and Ostriker 2012; Gatto et al. 2016; Howard et al. 2016). We adopt a sink particle model based on the methods of Federrath et al. (2010b) and Howard et al. (2016) to solve the star formation process in our work..
@@
These criteria are commonly used for sink particle implementations in mesh-based fluid dynamic solvers (e.g. Krumholz et al. 2004; Federrath et al. 2010b; Gong and Ostriker 2012). The criterion requiring the cell to be gravitationally bound has also been adopted in many studies. However, some studies indicate that stars can form in areas that are not gravitationally bound, such as the fast-expanding shell of an SNR bubble or a wind bubble (Cantat-Gaudin et al. 2018; Deharveng et al. 2009) or high-velocity fragments in the outflow area (Maiolino et al. 2017). To account for these scenarios, we relax the gravitational bound criterion in our simulations. The following subsections detail these criteria and the star formation process from sink particles.
@@ 2.3.2 Distance check
Even if gas clouds near a newly formed star cluster meet the density threshold for star formation, they can be disrupted by photoionization heating and radiative pressure from massive stars. In M82, star clusters are often embedded in molecular clouds with optical depths of up to $\tau\sim 10$. In such environments, radiative pressure may be comparable to supernovae in terms of momentum injection (Agertz et al. 2012). Combined with stellar winds, these early feedback processes can suppress star formation in neighboring regions. To account for this, we assign each sink particle a control volume that should not overlap with those of existing sink or star particles. The diameter of these spherical control volumes is set to three times the grid cell size.
@@ 2.3.3 Convergent flow
To confirm that a candidate cell is undergoing free-fall collapse, we check for convergent gas flow. This requires that the divergence of the velocity field , $\nabla\cdot v<0$, is negative in the central cell of the control volume (Gong and Ostriker 2012).
@@ 2.3.5 Gas accretion and star formation
Several methods exist for calculating gas accretion onto sink particles. Krumholz et al. (2004) categorize accretion into three scenarios based on the relative importance of pressure and gravity, as determined by the particle’s Bondi-Hoyle radius, and employ different accretion models accordingly. Gong and Ostriker (2012) introduce ‘ghost’ zones around particles and calculate accretion rates based on the flux at the zone boundaries. In our work, we adopt a simpler approach similar to Federrath et al. (2010b). Once a sink particle forms, gas accretion begins immediately. Grid cells within the control volume are continuously monitored, and any gas exceeding the density threshold $\rho_{th}$ is directly accreted by the sink particle. The mass accretion rate of a sink particle at each time step is given by:
@@
In our simulation, the radius of the control volume of a sink particle, $R_{cv}$, is $1.5$ times the size of the nearby grid cell. Since sink particles can form in regions with varying resolutions, their control volumes and the corresponding Jeans masses $M_{J}$ can differ. A fixed mass threshold, $M_{cnv}$, could lead to inconsistent distribution of stellar mass fraction, i.e. the ratio of stellar mass to the total mass in a star particle, across regions with different resolutions. To ensure a consistent distribution of the stellar mass fraction, we adopt a variable $M_{cnv}$. Based on the assumption that a) the total accreted mass is proportional to Jeans mass, b) sink particles of different sizes have a constant temperature, and c) the stellar mass fraction is independent of the particle mass, we derive $M_{cnv}\propto M_{J}\propto R_{cv}$. In our simulation, we adopt $M_{cnv}=5.0\times 10^{4}\,\rm{M_{\odot}}$ in regions have the finest grids with size $dx=4.0$ pc, and $M_{cnv}=1.0\times 10^{5}\,\rm{M_{\odot}}$ in regions with $dx=8.0$ pc, etc.
Under this star-formation scheme, the accreted gas is not fully converted into stars. Once the stellar-mass threshold is reached, the sink particle is converted into a star particle, which represents a stellar cluster containing both the newly formed stars and the remaining gravitationally bound gas. The detailed treatment of this residual gas is described in Section 2.4.1.
@@ 2.4 Feedback
Stellar feedback serves as the main driving force for the launch of the galactic outflow. Traditionally, the primary feedback mechanism considered in similar studies is the core-collapse SNe, which has been proven capable of launching galaxy-scale outflow in M82-like galaxies (e.g. Strickland and Stevens 2000; Cooper et al. 2008; Schneider et al. 2020). However, recent studies show that radiation feedback and stellar wind may play crucial roles in manipulating the effect of SNe feedback (Dale et al. 2014; Chevance et al. 2020). These early feedback processes may significantly lower the SFR and prevent the clustering of star formation. Some previous work suggests that efficient X-ray and extreme ultraviolet (EUV) radiation are capable of removing important coolants through photoionization, leading directly to the reduction of the accretion rate and SFR (e.g. Cantalupo 2009). Furthermore, stellar wind can create cavities and channels of ionized gas before the first SNe, mitigating the resistance an expanding SNR may meet. In this study, we consider three major feedback processes: radiation, stellar wind, and core-collapse SNe. We will examine how these mechanisms affect the starburst activity and the launch of outflow in M82. The effects of Type Ia SNe are ignored, as our simulations span only 30 Myr, during which their contribution to the feedback energy is expected to be minor compared to that of core-collapse supernovae. In our simulations, the stellar mass of a single star particle is equivalent to a medium-sized star cluster. Therefore, it can continuously produce stellar wind and radiation feedback and it will launch several individual SNe throughout its lifetime.
@@ 2.4.1 Gas return from sink particle
Most previous studies that adopted a sink particle-based star formation scheme have not included any returning mechanism for the remaining gas in sink particles. Some of these studies may assume that all the gas in sink particles will fully convert to stars (e.g. Wang et al. 2010; Padoan and Nordlund 2011; Gong and Ostriker 2012; Kim et al. 2018), which is valid only if the resolution can reach 0.1 parsec or higher. However, this treatment may overestimate the star formation efficiency (SFE and hereafter, i.e., the fraction of gas in GMCs/galaxy that has been converted into stars within a given time interval) in simulations with a resolution coarser than the size of the cores of GMCs. Meanwhile, other simulation studies on the evolution of GMC and ultra-faint dwarf galaxies believe that most unresolved gas in sink particles remains gravitationally bound to the cluster formed in GMC over a long time scale (Dale et al. 2014; Howard et al. 2016). Therefore, gas is stored in sink particles in these studies. However, in simulations with a relatively coarser resolution or with a long simulation time, such a treatment would cause a great fraction of the gas to be locked up in particles (Truelove et al. 1997; Hu et al. 2017), leading to an underestimation of the star formation efficiency. Overall, both approaches mentioned above have their own limitations. Recent studies indicate that early pre-supernovae feedback processes (i.e. stellar wind, radiation feedback) are capable of destroying the cloud and accelerating a great portion of the remaining gas to a gravitationally unbound state within 2 $\sim$ 15 Myr, in which the returning fraction may depend on the properties of the gas clumps (e.g., Kim et al. 2018; Li et al. 2019).
In our simulation, we use a simplified model to account for the dispersal of gas remaining in the sink/star particles. In the default run, we assume that all the remaining gas in the particles can eventually escape from their hosts. For comparison, we have also adopted two different setups in other runs, one with a gas return fraction of $50\%$, the other with the gas return fully disabled. The return of the remaining gas in a particle begins 2.0 Myr after star formation, based on previous simulations (e.g. Dale et al. 2014; Kim et al. 2018; Li et al. 2019; Fukushima and Yajima 2021). Moreover, the gas stored in a particle will return to the surrounding grid cells as an exponential function of time on a time scale $\tau_{rtn}$
$\dot{M}_{gas}=-\frac{M_{gas}}{\tau_{\rm{rtn}}}.$ (13)
Here in our simulation, a time scale of $\tau_{\rm{rtn}}=5.0Myr$ is adopted. At each time step, gas leaked from a particle is injected into grid cells within the particle’s control volume with an initial velocity equal to that of the host particle, and subsequently accelerated by stellar winds and radiation pressure applied to the same region.
@@ 2.4.2 Radiation
Multiple works have shown that radiation pressure and photoionization dominate over gas pressure in the central region of molecular clouds, acting as the main mechanisms for cloud destruction and disrupting star formation (Ceverino et al. 2013; Sales et al. 2013). More recently, radiation feedback has been proven capable of removing gas from the center of a molecular cloud and carving channels into the outer region, allowing supernova remnants to pass through without much resistance (Dale et al. 2014; Fujii et al. 2021). Emerick et al. (2018) shows that these two mechanisms can play a catalytic role in the launch of galactic outflow in a low-mass dwarf galaxy. To investigate the effect of radiation, we incorporate both radiation pressure and radiation heating feedback in our simulation. Note that photoionization is not explicitly included in our simulation. Instead, we employ photoelectric heating in our cooling/heating source term for the following reasons: 1) resolving the detailed chemical evolution is not the main intention of this work, and 2) most of the gas in M82 remains neutral throughout the starburst, while ionized gas in the dense area will recombine quickly. From a statistical point of view, photoionization can be treated as a heating term. In our simulation, only UV radiation is considered, as it carries the most energy and momentum at the early age of a starburst galaxy. We use two separate modules to account for the effect of the photoheating and radiation pressure separately, both of which will be described in this section.
@@
For heating rate calculation, we use the built-in photoelectric heating module of GRACKLE, which is based on Equation 1 in Wolfire et al. (1995). To compute photoelectric heating, the intensity of the interstellar radiation field must first be determined. To this end, we use an approach different from the optically thin assumption, considering that most parts of the M82’s gas disk, especially the star-forming clumps and SNR bubble shells, are opaque to UV radiation (Murray et al. 2009). Thompson and Quataert (2008) suggested that the characteristic opacity of UV radiation is $\kappa_{UV}\sim 10^{3}\,\rm{cm}^{2}\cdot\rm{g}^{-1}$ in the neutral gas. The two-temperature Planck mean opacity calculated from OPTAB (Hirose et al. 2021) also shows that for a typical OB star with an effective surface temperature of more than $10^{5}$ K surrounded by ambient gas with a temperature ranging from 10 to 20000 K, the opacity can be as high as $10^{2}\sim 10^{3}\,\rm{cm}^{2}\cdot\rm{g}^{-1}$. In such an environment, the inverse-square law could overestimate the intensity of the UV radiation field by a few orders of magnitude in a region far away from the UV source. To correct for this effect, we instead use a modified inverse-square law to estimate the intensity of the radiation field. More specifically, the UV flux distribution is obtained by solving the point-source divergence equation with absorption under spherical symmetry as follows.
@@
where $L$ is the length of the transmission path of the light. The IR optical depth $\tau_{\rm{IR}}$ in M82’s star cluster is assumed to be $10\sim 100$ based on the observed densities (Agertz et al. 2012), therefore IR may be the dominant term in radiation pressure. In the calculation of optical depth, we use a similar approach to that in the calculation of self-shielding in GRACKLE (Smith et al. 2016), we use the local Jeans length $R_{J}$ as the radius of the host clump within which the grid cell resides, therefore
@@
$\frac{dP}{dt}=-\nabla P_{\rm{Radiation}}$ (22)
@@ 2.4.3 Stellar Wind
In addition to radiation and SNe, the stellar wind is a non-negligible source of feedback, which could be comparable to that of SNe under certain circumstances (Agertz et al. 2012). We adopt a metallicity-dependent model proposed by Vink and Sander (2021) to calculate the stellar wind’s mass loss rate and final velocity.
Table 2: List of all the simulations performed, the first and second columns are the name and mass of gas disc, respectively. The third, fourth and fifth column indicate the setting of supernovae, stellar winds and radiation feedback, respectively. The sixth and seventh columns indicate the gas return fraction and initial gas metallicity respectively (see section 2.6 for more details). Simulation FD is specified as the fiducial run of our work. vSN (S16) stands for the variable SNe feedback model given by Sukhbold et al. (2016).
@@
GL00 $3.0\times 10^{9}\rm{M_{\odot}}$ vSN(S16) Yes Yes $\% 0.0$$02Z_{\odot}$